### v1

# Bridging Speech Emotion Recognition and Personality: Dataset and Temporal Interaction Condition Network

Abstract
This study investigates the interaction between personality traits and emotional expression, exploring how personality information can improve speech emotion recognition (SER). We collected personality annotation for the IEMOCAP dataset, and the statistical analysis identified significant correlations between personality traits and emotional expressions. To extract finegrained personality features, we propose a temporal interaction condition network (TICN), in which personality features are integrated with Hubert-based acoustic features for SER. Experiments show that incorporating ground-truth personality traits significantly enhances valence recognition, improving the concordance correlation coefficient (CCC) from 0.698 to 0.785 compared to the baseline without personality information. For practical applications in dialogue systems where personality information about the user is unavailable, we develop a front-end module of automatic personality recognition. Using these automatically predicted traits as inputs to our proposed TICN model, we achieve a CCC of 0.776 for valence recognition, representing an 11.17% relative improvement over the baseline. These findings confirm the effectiveness of personality-aware SER and provide a solid foundation for further exploration in personality-aware speech processing applications.
Index Terms:
Speech emotion recognition, Big Five personality traits, human computer interaction.
I Introduction
Speech emotion recognition (SER) is widely known as a vital component of natural human–computer interaction [1, 2]. Emotional information in speech reveals not only the immediate affective state of the speaker but also provides insight into how we think, feel, and behave [3]. By enabling intelligent systems to detect and respond to the emotional state of the user, SER makes interactions more intuitive and improves the overall user experience. Applications of SER range from empathetic speech assistants [4] and personalized systems for emotion-aware healthcare and automotive interfaces [5], underscoring its significance within the broader field of affective computing.
Moreover, how we express and manage emotions is significantly influenced by our personality traits [6]. These traits shape cognitive processes, reactions, and communication styles, and thus play a key role in social interactions and emotional expression. Among various personality models [7, 8], [9], the Big Five traits [10, 11] is the most widely adopted framework in psychological research [12, 13]. It defines five traits: openness (OP), conscientiousness (CO), extraversion (EX), agreeableness (AG), and neuroticism (NE). Each trait represents a distinct dimension of personality and offers a practical solution to quantify individual differences [14]. Previous studies have shown that these traits can substantially influence emotion perception, regulation, and expression [15, 16]. For instance, individuals high in EX tend to express their emotions more openly and with higher arousal, whereas those high in NE are often more sensitive to negative emotions and exhibit greater emotional instability.
In this work, we annotate the well-known Interactive Emotional Dyadic Motion Capture (IEMOCAP) dataset [17] with Big Five traits. This personality-annotated IEMOCAP (PA-IEMOCAP) dataset, to our knowledge, is the first dataset of this type for speech, enables direct investigation of the interplay between personality and emotion expression  [^1] The GitHub repository containing the data will be made publicly available when the paper is accepted.. We first examine the correlation between each Big Five trait and the emotional dimensions of valence and arousal. Additionally, we incorporate a straightforward linear layer to project personality traits as feature embedding, which is integrated with acoustic features to improve the performance of SER.
In practical applications, explicit personality information about users is unavailable [18]. In this work, we explore effective models that incorporate automatically predicted personality traits for improving SER. Our approach begins by establishing a baseline for personality recognition (PR) at the utterance level. We then conduct multi-task learning experiments, jointly training PR and SER to explore their mutual influence within a shared feature extractor. Given that personality traits remain consistent throughout a conversation, we further provide a conversational-level PR baseline. These experiments reveal the potential of real dialogue systems in personalizing response styles according to the personality of users.
Additionally, the predicted personality traits can serve as input of condition network, replacing ground-truth personality labels and thereby enable dialogue system better interpret and recognize emotional states about user. To better incorporate this predicted personality information for SER, we introduce a temporal interaction condition network (TICN). Although personality remains consistent over time, its impact on emotion expressions can vary across different speech segments. The proposed TICN is designed to capture this temporal variation, extracting finegrained personality features that enhance SER performance. Instead of combining features by concatenation, we incorporate cross attention mechanism [19], allowing effective interaction between personality and acoustic features.
The main contributions of our work are summarized as:
  • Personality annotated IEMOCAP dataset: We augment the widely-used IEMOCAP dataset with Big Five personality trait annotations, providing the first speech dataset that includes both emotion and personality labels.
  • Personality Recognition: We conduct PR experiments on PA-IEMOCAP dataset at both the utterance and conversation levels, establishing baseline performance for PR.
  • Personality-Conditioned SER Models: We propose TICN for incorporating annotated personality traits to enhance the SER model. We also evaluate whether leveraging predicted personality labels can improve SER performance when the personality information about the user is unavailable.
  • Multi-task Learning Frameworks: We explore multi-task learning frameworks for jointly modeling SER and PR, allowing us to investigate the mutual influence between these two tasks.
II Related Works
II-A Attribute Information for SER
SER has attracted increasing attention in recent years. Prior studies have shown that including additional speaker attributes such as gender [20], vocal characteristics [21], and linguistic content [22] can improve SER model performance. Researchers introduce multi-task learning to integrate these speaker attributes into SER models [23, 24], [25]. By simultaneously optimizing multiple related tasks, multi-task learning enables models to learn shared features and better utilize complementary information for emotion recognition [26, 27]. This approach has shown promising performance in SER [28, 29] benefiting from the extra information provided by emotion-related tasks.
Regarding the relationship between linguistic content and emotions, Cai et al. [22] proposed a multi-task learning method combining ASR and SER using the wav2vec 2.0 feature extractor. Joint training for ASR and SER allowed the model to leverage linguistic information effectively, leading to state-of-the-art performance on the IEMOCAP dataset. An ablation study further confirmed that carefully choosing the weight parameters for ASR significantly improved results. Additionally, speaker attributes have been widely studied: different speakers express emotions differently, affecting both acoustic and linguistic content for the same emotions. Thus, speaker-dependent SER systems usually outperform speaker-independent ones [30]. As a broad category of speakers, gender information can also improve performance in speaker-independent SER [20, 31]. More recently, Sharma et al. [28] proposed a multi-task learning approach using pretrained self-supervised learning feature extractor. Their study showed further advantages of multi-task learning by integrating additional tasks like language classification and regression tasks related to pitch and energy. These additional tasks improved SER performance significantly across 25 datasets covering 13 languages and 7 emotion categories, demonstrating the effectiveness of multi-task learning in enhancing model generalization, especially in multilingual settings. Despite the known relationship between personality traits and emotional expressions, personality remains an underexplored auxiliary attribute in SER.
II-B Joint Analysis of Personality and Emotion in Speech
Personality traits and emotional states are closely connected, both influencing how people express themselves [32, 33]. Understanding how these two aspects interact can lead to better performance in SER tasks. Recently, researchers have started exploring joint analysis frameworks to capture the interaction between personality and emotion, even though the lack of datasets annotated with both attributes remains a major challenge. Zhang et al. [34] proposed PersEmoN, a deep multi-task learning framework for jointly analyzing personality and emotion. Because no publicly available corpus contained both emotion and personality annotations, they used two separate datasets (one labeled for emotion and another for personality) and incorporated an adversarial learning method to mitigate the mismatch of different datasets. Their results highlighted an inherent relationship between personality and emotion: personality shows stable impact on how emotions are expressed and managed. For example, they observed that EX corresponds to more intense emotions, while NE correlates with more frequent negative expression. Similarly, Li et al. [35] explored the personality-emotion connection in a transfer learning manner. They trained a wav2vec2-based model [36] for SER and then finetuned it for PR, finding that emotional features, particularly the arousal dimension, significantly improve PR performance. These studies underscore the effectiveness of incorporating personality information into SER.
However, due to the lack of publicly available datasets that include annotations for both emotion and personality, previous works rely on separate data sources leading to substantial domain mismatch, and thus results in low generalizability of their findings.
III Personality Annotation for IEMOCAP dataset
We conducted personality annotations for the widely-used Interactive Emotional Dyadic Motion Capture (IEMOCAP) dataset to explore relationships between personality traits and emotional expressions. The IEMOCAP dataset [17] is a benchmark dataset extensively used in emotion recognition and sentiment analysis research. It comprises approximately 12 hours of multimodal data collected from 10 professional actors (5 males and 5 females), including high-quality audio recordings, video, detailed facial motion capture data, and textual transcriptions. Data collection involved five dyadic sessions, each including one male and one female actor performing scripted conversations and engaging in improvisational scenarios designed to elicit specific emotional expressions. The audio was captured using two microphones at a 48 kHz sampling rate and subsequently downsampled to 16 kHz to align with the common audio processing standard. Below, we introduce key aspects of our annotation method, consistency evaluation, and correlation analysis.
III-A Labeling Procedure
In the IEMOCAP dataset, professional actors perform diverse roles in each conversation. We presume they play different characters in each conversation, and we annotated the Big Five personality traits at the conversation level, resulting in 302 personality profiles for 151 dyadic interactions. The annotation was performed by six independent raters recruited via Amazon Mechanical Turk, each tasked with evaluating the personality traits of individual speakers based on both the audio recordings and transcribed text of the conversations. The Ten Item Personality Measure (TIPI) [37], a widely validated instrument for assessing the Big Five personality dimensions (OP, CO, EX, AG, and NE), was employed, with raters providing scores on a 7-point Likert scale (ranging from 1 = “strongly disagree” to 7 = “strongly agree”) for each trait. To ensure the quality and consistency of the annotations, we implemented a rigorous data-cleaning procedure. Specifically, we excluded ratings deemed contradictory, such as instances where a speaker was simultaneously rated as “critical, quarrelsome” and “sympathetic, warm” for the Agreeableness trait, as such inconsistencies could undermine the reliability of the personality profiles. Furthermore, to mitigate the impact of outliers and enhance robustness, we removed the single rating furthest from the median score for each trait, subsequently computing the average of the remaining five ratings to generate a final trait score.
TABLE I: Inter-annotator agreement analysis for Big Five personality trait annotations. We report both ICC(2,k) and Fleiss’ kappa values to evaluate the consistency among annotators.
Personality traits ICC Fleiss’s kappa
OP 0.89 0.49
CO 0.89 0.53
EX 0.89 0.47
AG 0.97 0.83
NE 0.95 0.71
Average 0.92 0.61
TABLE II: Pearson correlation coefficients (PCC) between the Big Five personality traits and arousal and valence. Values marked with indicate a strong correlation ($>|0.5|$).
Personality traits Valence Arousal
OP $0.53^{\dagger}$ $0.09$
CO $0.32$ $-0.22$
EX $0.35$ $0.32$
AG $0.51^{\dagger}$ $-0.11$
NE $-0.45$ $0.19$
III-B Consistency Assessment
To evaluate the reliability of personality annotations, we introduce two metrics: intraclass correlation coefficient (ICC) [38] for the collected continuous annotation (1-7 scale) and Fleiss’ kappa [39] for binarized (median-split) traits. As shown in Table I, ICC values demonstrate excellent agreement across all traits, ranging from 0.89 to 0.97 for the entire database, indicating strong reliability in the continuous personality assessments. We also ensure annotation reliability by converting the continuous personality trait scores into binary labels, maintaining robustness in simple personality classifications (e.g., extrovert vs. introvert). We observe that the average $\kappa$ across all annotations was 0.61. Among individual traits, AG and NE exhibited the highest agreement ($\kappa$=0.83 and $\kappa$=0.71, respectively), while EX had the lowest ($\kappa$=0.47). These results confirm the reliability and consistency of our personality annotations. The resultant annotated dataset is thereafter referred to as PA-IEMOCAP.
III-C Correlation of Big Five Traits and Emotions
We introduce Pearson correlation coefficients (PCC) [40] to assess the correlation between personality traits and the emotion dimensions of valence and arousal (Table II). Statistically significant correlations ($p<0.05$) were observed, notably OP showing strong positive correlations with valence (PCC = 0.53), which implies that individuals scoring high in OP tend to exhibit more positive emotions. NE, conversely, shows a pronounced negative correlation with valence (PCC = -0.45), indicating individuals with higher NE levels express more negative emotions. AG and CO also present noticeable associations with valence. Regarding arousal, EX has a positive correlation (PCC = 0.32), suggesting extroverts typically display heightened emotional energy. These results provide the first quantitative analysis of personality and emotion expression.
Figure 1: Overall flowchat of the proposed approach. We use all the utterances of a whole conversation to predict personality traits. Note that, we conduct independent experiments for prediction of each personality traits. The predicted traits are then projected by the proposed temporal interaction condition network (TICN) for improving SER.
IV SER with Ground-Truth Personality Traits
In the previous section, we confirmed a strong correlation between Big Five personality traits and emotion expression, particularly valence. Based on these findings, in this work, we explore the interplay between SER and PR in deep learning based models to validate the applications of these correlations in dialogue systems. First, we investigate whether incorporating personality information can enhance SER performance. Specifically, we introduce a condition network to project the Big Five personality traits into latent features. These personality features are then combined with acoustic features to assess whether personality information can benefit the model for emotion understanding.
Given an utterance $U\in\{U_{1},U_{2},...,U_{n}\}$, we incorporate HuBERT [41], which consists of 7 convolutional layers followed by 12 Transformer layers as acoustic feature extractor. This model is pretrained in a self-supervised learning manner and can capture both acoustic features and content information from the input utterance. The acoustic features learned from HuBERT are represented as $A^{e}\in\mathbb{R}^{t\times d}$, where $t$ is the number of temporal frames and $d$ is the hidden dimension of the Transformer layers. Previous works have validated that emotion perception requires both acoustic features and linguistic content. Therefore, we incorporate an ASR component to learn the linguistic information and thus ensure promising performance for SER [22, 42]. We feed $A^{e}$ into a linear layer and apply the connectionist temporal classification (CTC) loss function [43]:
$\mathcal{L}_{\text{ASR}}=\text{CTC}(y,\hat{y}),$ (1)
where $y$ represents the ground-truth transcription, and $\hat{y}$ denotes the predicted probability sqequence. Subsequently, we apply mean pooling across the time dimension of $A^{e}$ to obtain the utterance-level feature $A^{e}_{\text{pooled}}\in\mathbb{R}^{d}$, which is then used for either valence or arousal recognition:
$\mathcal{L}_{\text{SER}}=\left(e-\hat{e}\right)^{2},$ (2)
where the predicted emotion $\hat{e}$ is derived from $A^{e}_{\text{pooled}}$, and $e$ is the ground-truth emotion label. The overall loss function of the SER baseline integrates both ASR and SER objectives as follows:
$\mathcal{L}=(1-\lambda)\mathcal{L}_{\text{SER}}+\lambda\mathcal{L}_{\text{ASR}},$ (3)
where $\lambda$ is a weight parameter that balances the contribution of $\mathcal{L}_{\text{ASR}}$ and $\mathcal{L}_{\text{SER}}$.
Inspired by the correlation between personality and emotion expression in Table II, we first introduce a linear layer to project the personality features $P$ from each Big Five traits $t\in\mathbb{\{}{OP,CO,EX,AG,NE}\}$. The pooled acoustic features $A^{e}_{\text{pooled}}$ and the personality features $P$ are then concatenated to form a combined feature $E\in\mathbb{R}^{d+e}$ for SER. We also provide the comparison of using all personality features as additional information in the experiment section.
V Personality Recognition and SER with predicted Personality Traits
In real-world dialogue system applications, users do not typically provide explicit information about their personality traits. Therefore, instead of relying on ground-truth personality traits, we explore the use of predicted personality traits for improving SER. We first conduct comprehensive PR experiments using PA-IEMOCAP, evaluating recognition performance at both the utterance and conversation levels. Given the predicted personality, our framework enables dialogue systems to infer personality traits directly from speech input. Furthermore, we introduce a TICN to capture the temporal dynamics of personality impact on emotional expression and incorporate cross-attention for effective feature fusion.
V-A Utterance-level Personality Recognition
To align with SER, we first introduce utterance-level PR. First, we implement utterance-level PR by employing the HuBERT for acoustic feature extraction. Given the significant role of linguistic content in this task, we integrate an ASR component similar to our SER framework, and we extract the utterance-level feature $A^{p}_{pooled}$ through mean pooling for PR with the following loss function:
$\mathcal{L}_{\text{PR}}=\left(p-\hat{p}\right)^{2},$ (4)
where $p\in\mathbb{\{}{p^{OP},p^{CO},p^{EX},p^{AG},p^{NE}}\}$ is the ground-truth personality traits, and $\hat{p}$ represents the corresponding predicted personality traits. The overall loss function for our PR baseline combines both PR and ASR objectives:
$\mathcal{L}=(1-\gamma)\mathcal{L}_{\text{PR}}+\gamma\mathcal{L}_{\text{ASR}},$ (5)
where $\lambda$ is a weight parameter balancing the contribution of these two tasks.
Second, building upon our separate SER and PR frameworks, we explore multi-task learning to jointly train both tasks at the utterance-level to explore the mutual influence of these two tasks. Given an input utterance $U$, we extract acoustic features $A^{p}$ for three objectives: ASR, SER, and PR. For the ASR component, we maintain the same approach and apply the CTC loss function. After applying mean pooling across the time dimension to obtain $A^{p}_{\text{pooled}}\in\mathbb{R}^{d}$, we feed this utterance-level representation for both SER and PR. The overall loss function for our multi-task learning framework integrates all three objectives:
$\mathcal{L}=(1-\alpha-\beta)\mathcal{L}_{\text{PR}}+\alpha\mathcal{L}_{\text{SER}}+\beta\mathcal{L}_{\text{ASR}},$ (6)
where $\alpha$ and $\beta$ are weight parameters that balance the contribution of each task, subject to the constraint $0\leq\alpha,\beta\leq 1$ and $\alpha+\beta\leq 1$.
V-B Conversational-level Personality Recognition
We extend to conversation-level PR, acknowledging that personality traits become more apparent across longer conversations. For a conversation consisting of multiple utterances unit $\{U_{1},U_{2},...,U_{n}\}$, we extract utterance-level acoustic features and then concatenate them as conversational-level feature:
$C^{p}=Concat[A_{pooled1}^{p};A_{pooled2}^{p};...;A_{pooledn}^{p}],$ (7)
where $A_{pooledj}^{p}$ denotes the utterance-level acoustic features for the $j$-th utterance. The conversation-level representation $C$ captures the holistic personality manifestation across multiple utterances, enabling more robust PR in dialogue systems. The automatic PR component serves as a basis for our subsequent exploration of how predicted personality traits can enhance SER.
V-C Temporal Interaction Condition Network (TICN)
Self-attention is shown to be effective for SER as emotional expression varies at different temporal segments of an utterance. We hypothesize the impact of personality traits also varies temporally. To model this dynamic effect, we propose a TICN to extract personality traits into finegrained personality features. Specifically, we unsqueeze the personality feature $P$ projected from linear layer as $P^{t}\in\mathbb{R}^{t\times d}$ to align their dimensions with $A^{p}$. We integrate the temporal context information of the input utterance by performing element-wise addition between $A^{p}$ and $P$, which can be denoted as:
$P=0.9\cdot P\oplus 0.1\cdot A^{p}$ (8)
Finally, these fused features are passed through a Transformer layer to learn the interaction of personality and emotion expression over time. To effectively integrate the finegrained personality features with acoustic features, we apply a cross attention mechanism, allowing the model to adaptively capture relevant interactions between the two input features:
$E^{\mathrm{TICN}}=\text{softmax}\left((P^{t}W^{Q})(AW^{K})^{T}/\sqrt{d_{k}}\right)(AW^{V})$ (9)
This proposed approach enables $A^{p}$ to selectively incorporate information from $P^{t}$ that is most relevant to emotional expressions at each temporal segment.
As shown in the right sub-figure of Fig. 1, the outputs of the cross-attention layer $E^{\mathrm{TICN}}$ are then pooled to obtain fixed-length representations, which are subsequently used for SER. The overall loss function of the SER is the same as Eq. (3).
TABLE III: Effect of ground-truth Big Five traits in arousal and valence recognition. We use “*” to denote statistically significant improvement (p $<$ 0.05) from the baseline.
Model Valence (CCC) Arousal (CCC)
Baseline 0.698 0.711
OP CO EX AG NE Concat. CA TICN-CA Concat. CA TICN-CA
  0.737*   0.755*   0.782* 0.707 0.707 0.707
0.716   0.739*   0.751* 0.713 0.708 0.718
0.712   0.745*   0.740* 0.719 0.720 0.716
  0.733*   0.754*   0.760* 0.709 0.711 0.708
0.726   0.750*   0.768* 0.712 0.720 0.715
  0.746*   0.770* 0.785* 0.707 0.717 0.711
(a) Concat.
(b) CA
(c) TICN-CA
Figure 2: Hexbin density plot of valence recognition using different condition network integrating all ground-truth Big Five traits.
VI Experimental Evaluation
This section systematically examines the influence of personality information on SER. To comprehensively assess this relationship, we conducted three distinct experiments. The first experiment evaluates the effect of ground-truth personality traits in enhancing SER performance. This is achieved by integrating personality traits into SER systems through three distinct approaches. The second component investigates PR across multiple contexts. We first establish a baseline by performing utterance-level PR. Then, we implement multi-task learning to simultaneously optimize for SER and PR, examining the mutual effects between these tasks. Considering that personality traits typically remain consistent throughout a conversation, we expand our analysis to conversational-level PR, allowing us to better capture the temporal dynamics inherent in entire conversations. The third component explores practical applications by utilizing predicted personality traits to improve SER performance. This experiment adopts the same condition network for Big Five traits as those employed in the first experiment, ensuring consistency across our evaluations.
VI-A Implementation and Setup
We implemented the proposed model using PyTorch and the Huggingface Transformers repository [44]. Our experiments utilize HuBERT-base [41] as the acoustic feature extractor, which was pretrained on 60,000 hours of Libri-Light data [45] and comprises a convolutional feature extractor followed by 12 Transformer encoder layers with 768-dimensional hidden representations. During training, we strategically froze the CNN feature extractor and the first 6 Transformer layers while fine-tuning the remaining 6 layers, employing the AdamW optimizer with a learning rate of 5e-5, a mini-batch size of 2 with gradient accumulation steps of 8. The model was trained with dynamic batch padding to accommodate varying input lengths. For multi-task learning experiments, the weight parameters $\lambda$, $\gamma$, and $\alpha$, $\beta$ were set to 0.1.
We split the PA-IEMOCAP dataset into a training set (sessions 2–4), a validation set (14 conversations from session 1), and a test set (another 14 conversations from session 1), ensuring speaker-independent evaluation. We employed concordance correlation coefficient (CCC) as the primary metric for both PR and SER and Word Error Rate (WER) for the ASR, with model checkpoints saved based on the best CCC performance on the validation set.
VI-B Effect of Ground-Truth Personality Traits for SER
In this study, we explored three approaches to integrate personality information with acoustic features for enhancing SER. The first method, Concat, involves conditioning personality features through a linear layer and concatenating them with acoustic features. The second method, CA, employs a cross-attention mechanism to combine personality and acoustic features, facilitating effective interactions between the two modalities. The third method, TICN-CA, incorporates TICN to model the temporal dynamics of personality traits, which are then integrated with acoustic features via cross-attention (CA). We compare the proposed approaches with baseline model that uses only speech input. The results are presented in Table III.
We observe that concatenating acoustic features with personality embeddings significantly improved valence recognition, particularly when incorporating features learned from OP and AG. This aligns with the correlation analysis, where OP and AG exhibited the strongest correlations with valence. CA facilitates more effective interactions between personality and acoustic features, leading to noticeable improvements. The proposed TICN captures the temporal dynamics of personality traits’ effects on emotional expression, thereby significantly improving valence recognition. According to Table III, the best performance was achieved by using TICN to project all five traits collectively, which were then combined with acoustic features via cross-attention. This approach yielded a CCC of 0.791 for valence recognition. These findings suggest that integrating all personality traits provides a more robust improvement in valence recognition compared to utilizing any single trait individually. However, improvements in arousal recognition remained limited across all personality traits, indicating that personality information show more influential impact in valence rather than arousal in emotional expressions.
Fig. 2 shows Hexbin density plots for valence recognition, where darker hexagons indicate higher prediction density. Since sample sizes differ across valence labels, density comparisons should be made vertically along the true-label axis. For example, in plot (a), although samples at the lowest valence appear lighter overall than middle valence samples, they have darker hexagons closer to the diagonal (perfect prediction, shown in red). In contrast, samples with middle valence labels (around 3) show a relatively even vertical distribution, indicating poorer prediction accuracy. The TICN-CA approach (c) clearly provides better predictions than the concatenation model (a). Its regression line aligns more closely with the diagonal, and its data points are more densely clustered near the ideal prediction line compared to both (a) and (b). Notably, TICN-CA shows improved prediction clusters for valence values between 2 and 3, a range typically challenging for valence recognition systems, representing a significant advancement over the baseline model (a).
TABLE IV: Experimental results of personality recognition for each trait. Results are reported using concordance correlation coefficient (CCC). Utt: Utterance-level experiment. Conv: Conversational-level experiment.
Experimental settings Setting 1 Setting 2 Setting 3
Training Utt Utt Conv
Inference Utt Conv Conv
OP 0.441 0.492 0.695
CO 0.333 0.368 0.748
EX 0.437 0.467 0.809
AG 0.547 0.597 0.843
NE 0.438 0.464 0.793
Avg. 0.439 0.478 0.778
(a) Emotion recognition results
(b) Personality recognition results
Figure 3: Comparison between single-task and multi-task learning on personality and emotion recognition. Results are reported using concordance correlation coefficient (CCC).
VI-C Personality Recognition
VI-C1 Utterance-level Personality Recognition
We develop a baseline system for PR using the PA-IEMOCAP dataset. Two experimental settings were considered. In setting 1, utterance-level predictions were performed by extracting acoustic features from each utterance using HuBERT, followed by direct estimation of personality traits. In setting 2, a postprocessing step was introduced since personality traits remain consistent throughout a conversation. Specifically, predictions from all utterances belonging to the same speaker within a conversation were averaged to produce a single prediction per speaker. The CCC was then computed using these aggregated predictions.
The results of the PR experiments are presented in Table IV. We observe that AG gets the best recognition performance, and CO is the most difficult to predict. This indicates that individual utterances contain personality information, and achieved moderate CCC across all Big Five traits. Applying post-processing by averaging predictions across all utterances from the same speaker, we observed consistent improvements for all five traits. The CCC values increased by an average of 0.048, with the most significant improvement observed for AG and OP. This result validates that postprocessing mitigates the variability inherent in utterance-level experiments, and personality traits manifest more consistently across multiple utterances.
VI-C2 Conversational-Level Personality Recognition
Personality traits remain consistent throughout an entire conversation, although their expression may vary across individual utterances. Consequently, a conversational-level approach can yield more accurate predictions of the Big Five traits. By leveraging all utterances, this method allows the model to estimate the personality traits considering the whole conversation, and achieve the best performance. In this section, we conducted conversational-level experiments by concatenating the features extracted from HuBERT for all utterances of each speaker (setting 3).
As shown in the rightmost column of Table IV, this approach significantly outperformed utterance-level analysis, with CCC improvements exceeding 0.2 across all traits. AG achieved the best performance, with a CCC of 0.843, while OP exhibited the smallest improvement, achieving 0.695 on CCC. These results validate that predicting personality traits should take into account information from the entire conversation, as conversation-level features effectively capture contextual variations. The significant improvement between these two settings underscores the importance of broader conversational context and speaking patterns in revealing critical personality cues unavailable in single utterances.
VI-C3 Multi-task Learning for Speech Emotion Recognition and Personality Recognition
As a common practice of leveraging emotional-related information, we employ a simple multi-task learning for SER and PR to investigate the mutual influence between these two tasks. Specifically, we incorporate Hubert as feature encoder, and the output feature are used for both SER and PR. Since SER is inherently conducted at the utterance-level, we incorporate utterance-level acoustic features for both SER and PR. Previous experiments have validated that personality information can benefit SER, motivating us to explore how these tasks can benefit through joint training.
The experimental results are shown in Fig. 3. For valence recognition, multi-task learning consistently outperforms the baseline across all personality traits, with improvements ranging from 0.008 to 0.04. In contrast, arousal recognition performance was degraded from the baseline except for EX. For PR, no improvement is observed, except for CO with valence recognition. These results suggest that simple multi-task learning cannot effectively improve SER and PR.
VI-D Conditioning Predicted Personality Traits for SER
TABLE V: Valence recognition performance when conditioning on predicted or ground-truth personality traits using TICN-CA. We evaluate how using different Big Five traits (left) and different PR approaches (middle) affects valence recognition. Results are reported using concordance correlation coefficient (CCC). We use “*” to denote statistically significant improvement (p $<$ 0.05) from the baseline.
Incorporated Traits Predicted Traits (PR approach) ground-truth
OP CO EX AG NE Setting 1 Setting 2 Setting 3 Oracle
0.701   0.749*   0.756*   0.782*
0.688 0.707 0.728   0.751*
0.692 0.693 0.735   0.740*
0.710 0.729   0.761*   0.760*
0.705 0.723   0.762*   0.768*
0.695   0.754* 0.776* 0.785*
(a) Setting 1
(b) Setting 2
(c) Setting 3
Figure 4: Hexbin density plot of valence recognition using different condition networks integrating predicted personality traits. We compare three different settings estimating the personality traits.
The results of Table III and Fig. 3 (a) have shown that incorporating personality information can benefit valence recognition. In this section, we conduct experiments using predicted personality labels and more sophisticated TICN for enhancing SER. The predicted labels were obtained from the three different personality recognition settings described in the previous section. To integrate the personality information, we employed TICN-CA, which showed the best performance in Table III. Table V summarizes the results.
These results indicate that incorporating predicted personality traits improves valence recognition, with the conversational-level PR approach yielding the highest performance across all settings. When individual traits were incorporated, OP, AG, and NE led to the most notable improvements, consistent with earlier findings. The integration of all five traits predicted by conversational-level PR achieved the best results. In contrast, incorporating traits predicted from single utterance showed limited performance. These results further validate the effectiveness of personality-aware SER and highlight the advantages of leveraging conversational context for more reliable personality trait estimation.
Fig. 4 shows hexbin density plots of valence recognition results using the predicted personality traits from the three settings. Compared with setting 1, we observe notably improved distribution in both setting 2 and 3, even though using postprocessing at the inference stage leads to modest improvement in PR performance.
VII Conclusions
This study investigates the intrinsic relationship between personality traits and emotional expression in speech, especially exploring how personality information can be effectively leveraged to improve speech emotion recognition (SER). We annotated the well-known IEMOCAP emotional dataset with personality labels and identified strong correlations between traits and emotional expression, particularly valence. Then we present a novel approach for personality-aware SER.
We first established comprehensive baselines for PR at both utterance and conversation levels. The conversation-level models consistently outperformed utterance-level models by over 0.2 concordance correlation coefficient (CCC) across all five personality traits, reflecting the stability of personality characteristics over extended interactions. To capitalize on the strong correlation between personality traits and emotional expression, we propose the temporal interaction condition network with cross attention for feature fusion (TICN-CA). The proposed approach is designed to capture the dynamic impact of predicted personality traits on emotional expression across temporal segments of speech. Our experimental results demonstrate that integrating personality traits, whether ground-truth or predicted, significantly improves SER performance, particularly for valence. Using ground-truth personality traits, TICN improved the CCC from 0.698 to 0.785, while using predicted traits from conversation-level PR experiments as input achieved a comparable CCC of 0.776.
These findings validate the efficacy of personality-aware SER systems and highlight their potential for real-world applications where explicit personality data is unavailable. By demonstrating that predicted personality traits can substantively enhance emotion recognition, our approach paves the way for more personalized and emotionally intelligent human-computer interaction systems. This work not only addresses the gap in available emotion and personality annotated datasets but also establishes a foundation for future research into multimodal and context-aware affective computing.
VIII Acknowledgment
This work was supported by JST SPRING (JPMJSP2110), JST Moonshot R&\&&D (JPMJPS2011).

### v2

@@
 Abstract
-This study investigates the interaction between personality traits and emotional expression, exploring how personality information can improve speech emotion recognition (SER).
-We collected personality annotation for the IEMOCAP dataset, and the statistical analysis identified significant correlations between personality traits and emotional expressions.
+This study investigates the interaction between personality traits and emotion expression, exploring how personality information can improve speech emotion recognition (SER).
+We collect the personality annotation for the IEMOCAP dataset, making it the first speech dataset that contains both emotion and personality annotations (PA-IEMOCAP), and enabling direct integration of personality traits into SER.
+Statistical analysis on this dataset identified significant correlations between personality traits and emotional expressions.
 To extract finegrained personality features, we propose a temporal interaction condition network (TICN), in which personality features are integrated with HuBERT-based acoustic features for SER.
@@
 These findings confirm the effectiveness of personality-aware SER and provide a solid foundation for further exploration in personality-aware speech processing applications.
-Index Terms:
-Speech emotion recognition, Big Five personality traits, human computer interaction.
 I Introduction
-Speech emotion recognition (SER) is widely known as a vital component of natural human–computer interaction [1, 2].
-Emotional information in speech reveals not only the immediate affective state of the speaker but also provides insight into how we think, feel, and behave [3].
+Speech emotion recognition (SER) is widely known as a vital component of natural human–computer interaction [cowie2001emotion, ramakrishnan2013speech].
+Emotional information in speech reveals not only the immediate affective state of the speaker but also provides insight into how we think, feel, and behave [vernon2006thinking].
 By enabling intelligent systems to detect and respond to the emotional state of the user, SER makes interactions more intuitive and improves the overall user experience.
-Applications of SER range from empathetic speech assistants [4] and personalized systems for emotion-aware healthcare and automotive interfaces [5], underscoring its significance within the broader field of affective computing.
-Moreover, how we express and manage emotions is significantly influenced by our personality traits [6].
+Applications of SER range from empathetic speech assistants [chatterjee2021real] and personalized systems for emotion-aware healthcare and automotive interfaces [dhuheir2021emotion], underscoring its significance within the broader field of affective computing.
+Moreover, how we express and manage emotions is significantly influenced by our personality traits [stemmler2010personality].
 These traits shape cognitive processes, reactions, and communication styles, and thus play a key role in social interactions and emotional expression.
-Among various personality models [7, 8], [9], the Big Five traits [10, 11] is the most widely adopted framework in psychological research [12, 13].
+Among various personality models [cattell1992handbook, ashton2007empirical], [myers2010gifts], the Big Five traits [costa1999five, costa2008revised] is the most widely adopted framework in psychological research [azucar2018predicting, simha2020big].
 It defines five traits: openness (OP), conscientiousness (CO), extraversion (EX), agreeableness (AG), and neuroticism (NE).
-Each trait represents a distinct dimension of personality and offers a practical solution to quantify individual differences [14].
-Previous studies have shown that these traits can substantially influence emotion perception, regulation, and expression [15, 16].
+Each trait represents a distinct dimension of personality and offers a practical solution to quantify individual differences [langston1997beliefs].
+Previous studies have shown that these traits can substantially influence emotion perception, regulation, and expression [deniz2011investigation, curtis2015relationship].
 For instance, individuals high in EX tend to express their emotions more openly and with higher arousal, whereas those high in NE are often more sensitive to negative emotions and exhibit greater emotional instability.
-In this work, we annotate the well-known Interactive Emotional Dyadic Motion Capture (IEMOCAP) dataset [17] with Big Five traits.
-This personality-annotated IEMOCAP (PA-IEMOCAP) dataset, to our knowledge, is the first dataset of this type for speech, enables direct investigation of the interplay between personality and emotion expression  [^1] The GitHub repository containing the data will be made publicly available when the paper is accepted..
+In this work, we annotate the well-known Interactive Emotional Dyadic Motion Capture (IEMOCAP) dataset [busso2008iemocap] with Big Five traits.
+This personality-annotated IEMOCAP (PA-IEMOCAP) dataset, to our knowledge, is the first dataset of this type for speech, enables direct investigation of the interaction between personality and emotion expression [^1] Data available at https://github.com/Kyoto-University-Speech-and-Audio/PA-IEMOCAP.
 We first examine the correlation between each Big Five trait and the emotional dimensions of valence and arousal.
 Additionally, we incorporate a straightforward linear layer to project personality traits as feature embedding, which is integrated with acoustic features to improve the performance of SER.
-In practical applications, explicit personality information about users is unavailable [18].
+In practical applications, explicit personality information about users is unavailable [wu2017implicit].
 In this work, we explore effective models that incorporate automatically predicted personality traits for improving SER.
@@
 Given that personality traits remain consistent throughout a conversation, we further provide a conversational-level PR baseline.
-These experiments reveal the potential of real dialogue systems in personalizing response styles according to the personality of users.
+These experiments reveal the potential of real dialogue systems in personalizing response according to the personality of users.
 Additionally, the predicted personality traits can serve as input of condition network, replacing ground-truth personality labels and thereby enable dialogue system better interpret and recognize emotional states about user.
@@
 The proposed TICN is designed to capture this temporal variation, extracting finegrained personality features that enhance SER performance.
-Instead of combining features by concatenation, we incorporate cross attention mechanism [19], allowing effective interaction between personality and acoustic features.
+Instead of combining features by concatenation, we incorporate cross attention mechanism [vaswani2017attention], allowing effective interaction between personality and acoustic features.
 The main contributions of our work are summarized as:
-  • Personality annotated IEMOCAP dataset: We augment the widely-used IEMOCAP dataset with Big Five personality trait annotations, providing the first speech dataset that includes both emotion and personality labels.
+  • PA-IEMOCAP dataset: We augment the widely-used IEMOCAP dataset with Big Five personality trait annotations, providing the first speech dataset that includes both emotion and personality labels.
   • Personality Recognition: We conduct PR experiments on PA-IEMOCAP dataset at both the utterance and conversation levels, establishing baseline performance for PR.
@@
 II-A Attribute Information for SER
-SER has attracted increasing attention in recent years.
-Prior studies have shown that including additional speaker attributes such as gender [20], vocal characteristics [21], and linguistic content [22] can improve SER model performance.
-Researchers introduce multi-task learning to integrate these speaker attributes into SER models [23, 24], [25].
-By simultaneously optimizing multiple related tasks, multi-task learning enables models to learn shared features and better utilize complementary information for emotion recognition [26, 27].
-This approach has shown promising performance in SER [28, 29] benefiting from the extra information provided by emotion-related tasks.
-Regarding the relationship between linguistic content and emotions, Cai et al.
-[22] proposed a multi-task learning method combining ASR and SER using the wav2vec 2.0 feature extractor.
-Joint training for ASR and SER allowed the model to leverage linguistic information effectively, leading to state-of-the-art performance on the IEMOCAP dataset.
-An ablation study further confirmed that carefully choosing the weight parameters for ASR significantly improved results.
+SER has attracted increasing attention in recent years [khan2025joint].
+Prior studies have shown that including additional speaker attributes such as gender [li2019improved], vocal characteristics [triantafyllopoulos2021deep], and linguistic content [cai2021speech] can improve SER model performance.
+Researchers introduce multi-task learning to integrate these speaker attributes into SER models [collobert2008unified, standley2020tasks], [zhang2021survey].
+By simultaneously optimizing multiple related tasks, multi-task learning enables models to learn shared features and better utilize complementary information for emotion recognition [zhang2017cross, sener2018multi].
+This approach has shown promising performance in SER [sharma2022multi, latif2022multitask] benefiting from the extra information provided by emotion-related tasks.
+In multimodal emotion recognition, combining speech and text consistently outperforms speech-only systems, as linguistic information provides complementary semantic cues.
+Khan et al.
+[khan2025memocmt] introduced a cross-modal transformer that integrates HuBERT-based speech features with BERT-based text features, achieving state-of-the-art results on benchmark datasets.
+These findings highlight the importance of jointly modeling acoustic and linguistic cues for emotion recognition.
+When explicit transcripts are unavailable, automatic speech recognition (ASR) can serve as an auxiliary task to implicitly capture linguistic information from speech.
+Cai et al.
+[cai2021speech] proposed a multi-task learning framework combining ASR and SER with the wav2vec 2.0 feature extractor, where joint training effectively leveraged linguistic information and achieved state-of-the-art performance on IEMOCAP.
+Their ablation study further demonstrated that carefully weighting the ASR loss was critical to performance gains.
 Additionally, speaker attributes have been widely studied: different speakers express emotions differently, affecting both acoustic and linguistic content for the same emotions.
-Thus, speaker-dependent SER systems usually outperform speaker-independent ones [30].
-As a broad category of speakers, gender information can also improve performance in speaker-independent SER [20, 31].
+Thus, speaker-dependent SER systems usually outperform speaker-independent ones [gat2022speaker].
+As a broad category of speakers, gender information can also improve performance in speaker-independent SER [li2019improved, pappagari2020x].
 More recently, Sharma et al.
-[28] proposed a multi-task learning approach using pretrained self-supervised learning feature extractor.
+[sharma2022multi] proposed a multi-task learning approach using pretrained self-supervised learning feature extractor.
 Their study showed further advantages of multi-task learning by integrating additional tasks like language classification and regression tasks related to pitch and energy.
@@
 II-B Joint Analysis of Personality and Emotion in Speech
-Personality traits and emotional states are closely connected, both influencing how people express themselves [32, 33].
+Personality traits and emotional states are closely connected, both influencing how people express themselves [sagha2017effect, guidi2019analysis].
 Understanding how these two aspects interact can lead to better performance in SER tasks.
@@
 Zhang et al.
-[34] proposed PersEmoN, a deep multi-task learning framework for jointly analyzing personality and emotion.
+[zhang2019persemon] proposed PersEmoN, a deep multi-task learning framework for jointly analyzing personality and emotion.
 Because no publicly available corpus contained both emotion and personality annotations, they used two separate datasets (one labeled for emotion and another for personality) and incorporated an adversarial learning method to mitigate the mismatch of different datasets.
@@
 Similarly, Li et al.
-[35] explored the personality-emotion connection in a transfer learning manner.
-They trained a wav2vec2-based model [36] for SER and then finetuned it for PR, finding that emotional features, particularly the arousal dimension, significantly improve PR performance.
+[li2023transfer] explored the personality-emotion connection in a transfer learning manner.
+They trained a wav2vec2-based model [baevski2020wav2vec] for SER and then finetuned it for PR, finding that emotional features, particularly the arousal dimension, significantly improve PR performance.
 These studies underscore the effectiveness of incorporating personality information into SER.
@@
 III Personality Annotation for IEMOCAP dataset
-We conducted personality annotations for the widely-used Interactive Emotional Dyadic Motion Capture (IEMOCAP) dataset to explore relationships between personality traits and emotional expressions.
-The IEMOCAP dataset [17] is a benchmark dataset extensively used in emotion recognition and sentiment analysis research.
+We conducted personality annotations for the widely-used IEMOCAP dataset to explore relationships between personality traits and emotional expressions.
+The IEMOCAP dataset [busso2008iemocap] is a benchmark dataset extensively used in emotion recognition and sentiment analysis research.
 It comprises approximately 12 hours of multimodal data collected from 10 professional actors (5 males and 5 females), including high-quality audio recordings, video, detailed facial motion capture data, and textual transcriptions.
@@
 III-A Labeling Procedure
-In the IEMOCAP dataset, professional actors perform diverse roles in each conversation.
-We presume they play different characters in each conversation, and we annotated the Big Five personality traits at the conversation level, resulting in 302 personality profiles for 151 dyadic interactions.
+To elicit genuine emotional expressions and provide a more natural representation of affective behavior, the IEMOCAP dataset employs professional actors who perform diverse roles across conversations.
+We assume that the actors portrayed different characters in a manner closely approximating real-life behavior, and we therefore annotated the Big Five personality traits at the conversation level.
+These annotations reflected the enacted personalities in each conversation, rather than the actual personalities of the speakers, resulting in 302 personality profiles derived from 151 dyadic interactions.
 The annotation was performed by six independent raters recruited via Amazon Mechanical Turk, each tasked with evaluating the personality traits of individual speakers based on both the audio recordings and transcribed text of the conversations.
-The Ten Item Personality Measure (TIPI) [37], a widely validated instrument for assessing the Big Five personality dimensions (OP, CO, EX, AG, and NE), was employed, with raters providing scores on a 7-point Likert scale (ranging from 1 = “strongly disagree” to 7 = “strongly agree”) for each trait.
+To ensure annotation accuracy, annotators were recruited from the same sociocultural background as the original IEMOCAP speakers, i.e., native speakers of American English.
+The Ten Item Personality Measure (TIPI) [gosling2003very], a widely validated instrument for assessing the Big Five personality dimensions (OP, CO, EX, AG, and NE), was employed, with raters providing scores on a 7-point Likert scale (ranging from 1 = “strongly disagree” to 7 = “strongly agree”) for each trait.
 To ensure the quality and consistency of the annotations, we implemented a rigorous data-cleaning procedure.
-Specifically, we excluded ratings deemed contradictory, such as instances where a speaker was simultaneously rated as “critical, quarrelsome” and “sympathetic, warm” for the Agreeableness trait, as such inconsistencies could undermine the reliability of the personality profiles.
-Furthermore, to mitigate the impact of outliers and enhance robustness, we removed the single rating furthest from the median score for each trait, subsequently computing the average of the remaining five ratings to generate a final trait score.
-TABLE I: Inter-annotator agreement analysis for Big Five personality trait annotations. We report both ICC(2,k) and Fleiss’ kappa values to evaluate the consistency among annotators.
-Personality traits ICC Fleiss’s kappa
-OP 0.89 0.49
-CO 0.89 0.53
-EX 0.89 0.47
-AG 0.97 0.83
-NE 0.95 0.71
-Average 0.92 0.61
-TABLE II: Pearson correlation coefficients (PCC) between the Big Five personality traits and arousal and valence. Values marked with indicate a strong correlation ($>|0.5|$).
-Personality traits Valence Arousal
-OP $0.53^{\dagger}$ $0.09$
-CO $0.32$ $-0.22$
-EX $0.35$ $0.32$
-AG $0.51^{\dagger}$ $-0.11$
-NE $-0.45$ $0.19$
-III-B Consistency Assessment
-To evaluate the reliability of personality annotations, we introduce two metrics: intraclass correlation coefficient (ICC) [38] for the collected continuous annotation (1-7 scale) and Fleiss’ kappa [39] for binarized (median-split) traits.
-As shown in Table I, ICC values demonstrate excellent agreement across all traits, ranging from 0.89 to 0.97 for the entire database, indicating strong reliability in the continuous personality assessments.
-We also ensure annotation reliability by converting the continuous personality trait scores into binary labels, maintaining robustness in simple personality classifications (e.g., extrovert vs.
+Specifically, we excluded ratings deemed contradictory, such as instances where a speaker was simultaneously rated as “critical, quarrelsome” and “sympathetic, warm” for the AG trait, as such inconsistencies could undermine the reliability of the personality profiles.
+All problematic annotations were discarded, and new raters were recruited to replace them, ensuring that each conversation ultimately received six valid and independent ratings in total.
+The final trait score was computed as the average of the six ratings, without applying any further normalization.
+TABLE I: Inter-annotator agreement analysis for Big Five personality trait annotations.
+We report ICC(2,k) with 95% confidence intervals (CI 95%) and Fleiss’ kappa to evaluate the consistency among annotators.
+ICC(3,k) is additionally reported as a sensitivity check.
+Personality traits ICC(2,k), CI 95% ICC(3,k) Fleiss’ kappa
+OP 0.89, [0.87, 0.91] 0.90 0.49
+CO 0.89, [0.87, 0.91] 0.89 0.53
+EX 0.89, [0.87, 0.91] 0.89 0.47
+AG 0.97, [0.96, 0.97] 0.97 0.83
+NE 0.95, [0.94, 0.96] 0.95 0.71
+Average 0.92 0.92 0.61
+TABLE II: Pearson correlation coefficients (PCC) between the Big Five personality traits and arousal and valence at the utterance level.
+Subgroup analyses are reported for male and female speakers.
+Values marked with indicate a strong correlation ($|PCC|>0.5$).
+Overall ($N=10039$) Male ($N=5239$) Female ($N=4800$)
+Personality Valence Arousal Valence Arousal Valence Arousal
+[…]
+CO $0.33$ 0.21-0.21 $0.28$ 0.19-0.19 $0.38$ 0.24-0.24
+EX $0.35$ $0.32$ $0.35$ $0.29$ $0.36$ $0.34$
+[…]
+[…]
+TABLE III: Pearson correlation coefficients (PCC) between the Big Five personality traits and arousal and valence at the conversation level.
+Subgroup analyses are reported for male and female speakers.
+Values marked with indicate a strong correlation ($|PCC|>0.5$).
+Overall ($N=302$) Male ($N=151$) Female ($N=151$)
+Personality Valence Arousal Valence Arousal Valence Arousal
+[…]
+CO $0.41$ 0.41-0.41 $0.33$ 0.36-0.36 $0.49$ 0.48-0.48
+[…]
+[…]
+[…]
+To evaluate the reliability of personality annotations, we employed two metrics: the intraclass correlation coefficient (ICC) [fisher1970statistical] for the continuous ratings (1–7 scale) and Fleiss’ kappa [fleiss2013statistical] for the binarized traits obtained via median split.
+Before the consistency analysis, we excluded the rating that deviated the most from the median for each trait.
+As shown in Table I, inter-annotator agreement was excellent across all traits, with an average ICC(2,k) of 0.92.
+After data cleaning, the ICC(2,k) values ranged from 0.89 to 0.97, with 95% confidence intervals consistently above 0.85.
+As a sensitivity check, we additionally report ICC(3,k), which yielded nearly identical values, confirming the robustness of the reliability estimates.
+Without trimming the ratings furthest from the median score (mentioned in Section 3(A)), the average ICC(2,k) was 0.83, which still indicates strong reliability.
+We also ensure annotation reliability by converting the continuous personality trait scores into binary labels using a median split, thereby maintaining robustness in simplified personality classifications (e.g., extrovert vs.
 introvert).
-We observe that the average $\kappa$ across all annotations was 0.61.
-Among individual traits, AG and NE exhibited the highest agreement ($\kappa$=0.83 and $\kappa$=0.71, respectively), while EX had the lowest ($\kappa$=0.47).
+The average Fleiss’ $\kappa$ across all traits was 0.61; without trimming extreme ratings, the average $\kappa$ decreased to 0.39.
+Among individual traits, AG and NE exhibited the highest agreement ($\kappa$=0.83 and $\kappa$=0.71, respectively), while EX had the lowest ($\kappa$=0.47), likely because the conversational context did not clearly reflect introversion–extroversion, making it harder to annotate.
 These results confirm the reliability and consistency of our personality annotations.
 The resultant annotated dataset is thereafter referred to as PA-IEMOCAP.
-III-C Correlation of Big Five Traits and Emotions
-We introduce Pearson correlation coefficients (PCC) [40] to assess the correlation between personality traits and the emotion dimensions of valence and arousal (Table II).
-Statistically significant correlations ($p<0.05$) were observed, notably OP showing strong positive correlations with valence (PCC = 0.53), which implies that individuals scoring high in OP tend to exhibit more positive emotions.
-NE, conversely, shows a pronounced negative correlation with valence (PCC = -0.45), indicating individuals with higher NE levels express more negative emotions.
-AG and CO also present noticeable associations with valence.
+III-B Correlation of Big Five Traits and Emotions
+We introduce Pearson correlation coefficients (PCC) [cohen2009pearson] between the Big Five traits and the two emotion dimensions, valence and arousal, at utterance level (Table II).
+Statistically significant correlations ($p<0.05$) were observed: OP shows strong positive correlations with valence (PCC = 0.53), which implies that individuals scoring high in OP tend to exhibit more positive emotions.
+NE shows a pronounced negative correlation with valence (PCC = -0.45), indicating individuals with higher NE levels express more negative emotions.
+AG and CO also exhibit noticeable associations with valence.
 Regarding arousal, EX has a positive correlation (PCC = 0.32), suggesting extroverts typically display heightened emotional energy.
-These results provide the first quantitative analysis of personality and emotion expression.
+Because personality labels are constant within a conversation and thus repeated across a speaker’s utterances, treating utterances as independent may violate the independence assumption.
+We therefore averaged valence and arousal for each speaker in a conversation and computed PCCs across conversations (Table III), thereby capturing the association between personality and overall emotion expression.
+Compared with the utterance-level analysis, correlations show modest increases overall; in particular, the association between Extraversion and arousal increases to (PCC=0.61).
+We hypothesize that aggregating at the speaker level reduces within-speaker noise, thereby yielding better correlations.
+In addition, we conducted subgroup analyses by gender.
+The results for male and female speakers show broadly consistent patterns with the overall population, and no substantial differences between the genders were observed.
+For instance, the OP–valence and EX–arousal associations remained strong across both groups, while NE consistently exhibited negative correlations with valence.
 Figure 1: Overall flowchat of the proposed approach.
@@
 These personality features are then combined with acoustic features to assess whether personality information can benefit the model for emotion understanding.
-Given an utterance $U\in\{U_{1},U_{2},...,U_{n}\}$, we incorporate HuBERT [41], which consists of 7 convolutional layers followed by 12 Transformer layers as acoustic feature extractor.
+Given an utterance $U\in\{U_{1},U_{2},...,U_{n}\}$, we incorporate HuBERT [hsu2021hubert], which consists of 7 convolutional layers followed by 12 Transformer layers as acoustic feature extractor.
 This model is pretrained in a self-supervised learning manner and can capture both acoustic features and content information from the input utterance.
@@
 Previous works have validated that emotion perception requires both acoustic features and linguistic content.
-Therefore, we incorporate an ASR component to learn the linguistic information and thus ensure promising performance for SER [22, 42].
-We feed $A^{e}$ into a linear layer and apply the connectionist temporal classification (CTC) loss function [43]:
-$\mathcal{L}_{\text{ASR}}=\text{CTC}(y,\hat{y}),$ (1)
+Therefore, we incorporate an ASR component to learn the linguistic information and thus ensure promising performance for SER [cai2021speech, gao2024enhancing].
+We feed $A^{e}$ into a linear layer and apply the connectionist temporal classification (CTC) loss function [graves2006connectionist]:
+[…]
 where $y$ represents the ground-truth transcription, and $\hat{y}$ denotes the predicted probability sqequence.
@@
 The overall loss function of the SER baseline integrates both ASR and SER objectives as follows:
-$\mathcal{L}=(1-\lambda)\mathcal{L}_{\text{SER}}+\lambda\mathcal{L}_{\text{ASR}},$ (3)
+$\mathcal{L}_{\text{SER\_baseline}}=(1-\lambda)\mathcal{L}_{\text{SER}}+\lambda\mathcal{L}_{\text{ASR}},$ (3)
 where $\lambda$ is a weight parameter that balances the contribution of $\mathcal{L}_{\text{ASR}}$ and $\mathcal{L}_{\text{SER}}$.
-Inspired by the correlation between personality and emotion expression in Table II, we first introduce a linear layer to project the personality features $P$ from each Big Five traits $t\in\mathbb{\{}{OP,CO,EX,AG,NE}\}$.
+Inspired by the correlation between personality and emotion expression in Table II and III, we first introduce a linear layer to project the personality features $P$ from each Big Five traits $t\in\mathbb{\{}{OP,CO,EX,AG,NE}\}$.
 The pooled acoustic features $A^{e}_{\text{pooled}}$ and the personality features $P$ are then concatenated to form a combined feature $E\in\mathbb{R}^{d+e}$ for SER.
@@
 Given the predicted personality, our framework enables dialogue systems to infer personality traits directly from speech input.
-Furthermore, we introduce a TICN to capture the temporal dynamics of personality impact on emotional expression and incorporate cross-attention for effective feature fusion.
+Furthermore, we introduce a TICN to capture the temporal dynamics of personality impact on emotional expression and incorporate cross-attention for effective feature fusion (Fig.
+1).
 V-A Utterance-level Personality Recognition
@@
 The overall loss function for our PR baseline combines both PR and ASR objectives:
-$\mathcal{L}=(1-\gamma)\mathcal{L}_{\text{PR}}+\gamma\mathcal{L}_{\text{ASR}},$ (5)
-where $\lambda$ is a weight parameter balancing the contribution of these two tasks.
+$\mathcal{L}_{\text{PR\_baseline}}=(1-\gamma)\mathcal{L}_{\text{PR}}+\gamma\mathcal{L}_{\text{ASR}},$ (5)
+where $\gamma$ is a weight parameter balancing the contribution of these two tasks.
 Second, building upon our separate SER and PR frameworks, we explore multi-task learning to jointly train both tasks at the utterance-level to explore the mutual influence of these two tasks.
@@
 The overall loss function for our multi-task learning framework integrates all three objectives:
-$\mathcal{L}=(1-\alpha-\beta)\mathcal{L}_{\text{PR}}+\alpha\mathcal{L}_{\text{SER}}+\beta\mathcal{L}_{\text{ASR}},$ (6)
+$\mathcal{L}_{\text{multitask}}=(1-\alpha-\beta)\mathcal{L}_{\text{PR}}+\alpha\mathcal{L}_{\text{SER}}+\beta\mathcal{L}_{\text{ASR}},$ (6)
 where $\alpha$ and $\beta$ are weight parameters that balance the contribution of each task, subject to the constraint $0\leq\alpha,\beta\leq 1$ and $\alpha+\beta\leq 1$.
@@
 Self-attention is shown to be effective for SER as emotional expression varies at different temporal segments of an utterance.
-We hypothesize the impact of personality traits also varies temporally.
-To model this dynamic effect, we propose a TICN to extract personality traits into finegrained personality features.
-Specifically, we unsqueeze the personality feature $P$ projected from linear layer as $P^{t}\in\mathbb{R}^{t\times d}$ to align their dimensions with $A^{p}$.
-We integrate the temporal context information of the input utterance by performing element-wise addition between $A^{p}$ and $P$, which can be denoted as:
-$P=0.9\cdot P\oplus 0.1\cdot A^{p}$ (8)
+Although the personality of a speaker remains consistent, we hypothesize that its influence on emotion expression varies across different temporal segments of an utterance.
+To enable the model to capture this dynamic alignment, we replicate $P$ across time steps, not to imply temporal variability in personality itself, but to allow personality information to interact with acoustic features at each segment.
+Through TICN, this design facilitates the extraction of fine-grained personality-conditioned features that reflect the varying salience of personality traits in shaping emotional expression over time.
+Specifically, we unsqueeze the personality feature $P$ projected from linear layer as $P^{t}\in\mathbb{R}^{t\times d}$ to align their dimensions with $A^{e}$ (acoustic feature for SER).
+We integrate the temporal emotion information of the input utterance by performing element-wise addition between $A^{e}$ and $P$, which can be denoted as:
+$P=0.9\cdot P\oplus 0.1\cdot A^{e}$ (8)
 Finally, these fused features are passed through a Transformer layer to learn the interaction of personality and emotion expression over time.
 To effectively integrate the finegrained personality features with acoustic features, we apply a cross attention mechanism, allowing the model to adaptively capture relevant interactions between the two input features:
-$E^{\mathrm{TICN}}=\text{softmax}\left((P^{t}W^{Q})(AW^{K})^{T}/\sqrt{d_{k}}\right)(AW^{V})$ (9)
-This proposed approach enables $A^{p}$ to selectively incorporate information from $P^{t}$ that is most relevant to emotional expressions at each temporal segment.
+$E^{\mathrm{ticn}}=\text{softmax}\left((P^{t}W^{Q})(A^{e}W^{K})^{T}/\sqrt{d_{k}}\right)(A^{e}W^{V})$ (9)
+Here, $P^{t}$ serve as the query, while $A^{e}$ act as both the key and the value, with the number of attention heads set to 4.
+We also examined the reverse configuration, in which $P^{t}$ acted as key/value and $A^{e}$ as query; however, this design yielded unsatisfactory results and was not further considered.
+This proposed approach enables $A^{e}$ to selectively incorporate information from $P^{t}$ that is most relevant to emotional expressions at each temporal segment.
 As shown in the right sub-figure of Fig.
-1, the outputs of the cross-attention layer $E^{\mathrm{TICN}}$ are then pooled to obtain fixed-length representations, which are subsequently used for SER.
-The overall loss function of the SER is the same as Eq.
+1, the outputs of the cross-attention layer $E^{\mathrm{ticn}}\in\mathbb{R}^{t\times d}$ are then pooled to obtain fixed-length representations, which are subsequently used for SER.
+The overall loss function of SER follows the same formulation as in Eq.
 (3).
-TABLE III: Effect of ground-truth Big Five traits in arousal and valence recognition. We use “*” to denote statistically significant improvement (p $<$ 0.05) from the baseline.
+TABLE IV: Effect of ground-truth Big Five traits in arousal and valence recognition. We use “*” to denote statistically significant improvement (p << 0.05) from the baseline.
 Model Valence (CCC) Arousal (CCC)
@@
 VI-A Implementation and Setup
-We implemented the proposed model using PyTorch and the Huggingface Transformers repository [44].
-Our experiments utilize HuBERT-base [41] as the acoustic feature extractor, which was pretrained on 60,000 hours of Libri-Light data [45] and comprises a convolutional feature extractor followed by 12 Transformer encoder layers with 768-dimensional hidden representations.
-During training, we strategically froze the CNN feature extractor and the first 6 Transformer layers while fine-tuning the remaining 6 layers, employing the AdamW optimizer with a learning rate of 5e-5, a mini-batch size of 2 with gradient accumulation steps of 8.
+We implemented the proposed model using PyTorch and the Huggingface Transformers repository [wolf2020Transformers].
+Our experiments utilize HuBERT-base [hsu2021hubert] as the acoustic feature extractor, which was pretrained on 60,000 hours of Libri-Light data [kahn2020libri] and comprises a convolutional feature extractor followed by 12 Transformer encoder layers with 768-dimensional hidden representations.
+For the linear layers in TICN, the output dimension was fixed at 768 to ensure consistency between $P^{t}$ and $A^{e}$.
+During training for both SER and PR tasks, the CNN feature extractor together with the first six Transformer layers were frozen, while the remaining six layers were fine-tuned.
+Optimization was performed using AdamW with a learning rate of $5\times 10^{-5}$, a mini-batch size of 2, and gradient accumulation over 8 mini-batches.
 The model was trained with dynamic batch padding to accommodate varying input lengths.
-For multi-task learning experiments, the weight parameters $\lambda$, $\gamma$, and $\alpha$, $\beta$ were set to 0.1.
-We split the PA-IEMOCAP dataset into a training set (sessions 2–4), a validation set (14 conversations from session 1), and a test set (another 14 conversations from session 1), ensuring speaker-independent evaluation.
-We employed concordance correlation coefficient (CCC) as the primary metric for both PR and SER and Word Error Rate (WER) for the ASR, with model checkpoints saved based on the best CCC performance on the validation set.
+For multi-task learning experiments, the weight parameters $\lambda$, $\gamma$, $\alpha$, and $\beta$ were tuned among 1,0.1,0.01{1,0.1,0.01}, with 0.1 yielding the best performance for all of them and thus adopted in our experiments.
+Each experiment was repeated three times with the default random seed of Huggingface and we report the average results.
+We split the PA-IEMOCAP dataset into a training set (sessions 2–4), a validation set (14 conversations from session 1), and a test set (another 14 conversations from session 1), ensuring speaker-independent evaluation for both SER and PR.
+We employed concordance correlation coefficient (CCC) [lawrence1989concordance] as the primary metric for both PR and SER, with model checkpoints saved based on the best CCC performance (SER) on the validation set.
 VI-B Effect of Ground-Truth Personality Traits for SER
@@
 We compare the proposed approaches with baseline model that uses only speech input.
-The results are presented in Table III.
+The results are presented in Table IV.
 We observe that concatenating acoustic features with personality embeddings significantly improved valence recognition, particularly when incorporating features learned from OP and AG.
@@
 The proposed TICN captures the temporal dynamics of personality traits’ effects on emotional expression, thereby significantly improving valence recognition.
-According to Table III, the best performance was achieved by using TICN to project all five traits collectively, which were then combined with acoustic features via cross-attention.
+According to Table IV, the best performance was achieved by using TICN to project all five traits collectively, which were then combined with acoustic features via cross-attention.
 This approach yielded a CCC of 0.791 for valence recognition.
+The proposed model introduces only a few additional linear layers, one Transformer block, and a cross-attention mechanism, resulting in a relatively small parameter increase of about 12.1% (106.8M vs.
+95.4M in the baseline).
 These findings suggest that integrating all personality traits provides a more robust improvement in valence recognition compared to utilizing any single trait individually.
@@
 Notably, TICN-CA shows improved prediction clusters for valence values between 2 and 3, a range typically challenging for valence recognition systems, representing a significant advancement over the baseline model (a).
-TABLE IV: Experimental results of personality recognition for each trait.
+TABLE V: Experimental results of personality recognition for each trait.
 Results are reported using concordance correlation coefficient (CCC).
@@
 The CCC was then computed using these aggregated predictions.
-The results of the PR experiments are presented in Table IV.
+The results of the PR experiments are presented in Table V.
 We observe that AG gets the best recognition performance, and CO is the most difficult to predict.
@@
 In this section, we conducted conversational-level experiments by concatenating the features extracted from HuBERT for all utterances of each speaker (setting 3).
-As shown in the rightmost column of Table IV, this approach significantly outperformed utterance-level analysis, with CCC improvements exceeding 0.2 across all traits.
-AG achieved the best performance, with a CCC of 0.843, while OP exhibited the smallest improvement, achieving 0.695 on CCC.
+As shown in the rightmost column of Table V, this approach significantly outperformed utterance-level analysis, with CCC improvements exceeding 0.2 across all traits.
+AG achieves the best performance, with a CCC of 0.843, while OP exhibites the smallest improvement, achieving 0.695 on CCC.
 These results validate that predicting personality traits should take into account information from the entire conversation, as conversation-level features effectively capture contextual variations.
-The significant improvement between these two settings underscores the importance of broader conversational context and speaking patterns in revealing critical personality cues unavailable in single utterances.
+The significant improvement between these two settings underscores the importance of broader conversational context and speaking patterns, suggesting that, much like human perception, computational models also require sufficient conversational evidence to form a reliable impression of a speaker’s personality.
 VI-C3 Multi-task Learning for Speech Emotion Recognition and Personality Recognition
@@
 VI-D Conditioning Predicted Personality Traits for SER
-TABLE V: Valence recognition performance when conditioning on predicted or ground-truth personality traits using TICN-CA.
+TABLE VI: Valence recognition performance when conditioning on predicted or ground-truth personality traits using TICN-CA.
 We evaluate how using different Big Five traits (left) and different PR approaches (middle) affects valence recognition.
 Results are reported using concordance correlation coefficient (CCC).
-We use “*” to denote statistically significant improvement (p $<$ 0.05) from the baseline.
+We use “*” to denote statistically significant improvement (p << 0.05) from the baseline.
 Incorporated Traits Predicted Traits (PR approach) ground-truth
@@
 Figure 4: Hexbin density plot of valence recognition using different condition networks integrating predicted personality traits. We compare three different settings estimating the personality traits.
-The results of Table III and Fig.
+The results of Table IV and Fig.
 3 (a) have shown that incorporating personality information can benefit valence recognition.
@@
 The predicted labels were obtained from the three different personality recognition settings described in the previous section.
-To integrate the personality information, we employed TICN-CA, which showed the best performance in Table III.
-Table V summarizes the results.
+To integrate the personality information, we employed TICN-CA, which showed the best performance in Table IV.
+Table VI summarizes the results.
 These results indicate that incorporating predicted personality traits improves valence recognition, with the conversational-level PR approach yielding the highest performance across all settings.
@@
 These results further validate the effectiveness of personality-aware SER and highlight the advantages of leveraging conversational context for more reliable personality trait estimation.
+Notably, the introduction of the conversational PR module increases the parameter count substantially (from 95.4M in the baseline SER model to 333.9M in total), which should be considered as the trade-off for the performance gain.
 Fig.
@@
 Compared with setting 1, we observe notably improved distribution in both setting 2 and 3, even though using postprocessing at the inference stage leads to modest improvement in PR performance.
+TABLE VII: Different levels of Gaussian noise injected into the personality labels during inference. Results are reported using concordance correlation coefficient (CCC).
+Gaussian noise Valence (CCC)
+0.0 (GT) 0.785
+0.1 0.782
+0.2 0.780
+0.4 0.778
+0.8 0.743
+1.2 0.697
+1.6 0.657
+In practical applications, a valence recognition model may be trained with ground-truth personality information but must rely on automatically predicted traits during inference, whose accuracy gradually improves as the dialogue progresses.
+To examine robustness under such conditions, we injected zero-mean Gaussian noise with varying standard deviations into the personality labels at inference and evaluated its impact on valence recognition.
+As shown in Table VII, the model remains stable when the noise standard deviation is below 0.8, achieving results comparable to those obtained with ground-truth personality labels (0.785 vs.
+0.743).
+This finding indicates that even coarse or approximate personality estimates are sufficient to provide meaningful benefits for valence recognition.
+Performance declines more noticeably under larger perturbations (e.g., CCC = 0.657 at $\sigma=1.6$), yet the results demonstrate that the model maintains considerable robustness to moderate levels of personality recognition error.
+VI-E Comparasion with existing approaches
+To facilitate comparison with prior studies, we also report the SER results under the leave-one-session-out (LOSO) data split.
+TABLE VIII: Comparison with existing valence recognition approaches.
+Approach Input Year CCC
+Srinivasan [srinivasan2022representation] Speech 2022 0.582
+Wagner [wagner2023dawn] Speech 2023 0.478
+Ispas [ispas2023multi] Speech & Text 2023 0.744
+Vlasenko [vlasenko2024comparing] Speech 2024 0.683
+Zhou [zhou2024learning] Speech 2024 0.674
+Messaoudi [messaoudi2024modeling] Text 2024 0.724
+Baseline (w/o traits) Speech 2025 0.681
+Proposed (Predicted traits) Speech 2025 0.736
+Proposed (GT traits) Speech 2025 0.766
+The results in Table VIII demonstrate the rapid progress of valence recognition systems over recent years.
+Notably, text-based models (e.g., [messaoudi2024modeling]) generally outperform speech-based counterparts, suggesting that lexical cues provide highly informative signals for valence estimation.
+Furthermore, multimodal approaches that jointly leverage speech and text (e.g., [ispas2023multi]) have achieved the highest reported CCC scores in prior studies, highlighting the complementary nature of linguistic and acoustic information in capturing valence.
+As shown in Table VIII, our baseline system achieved performance comparable to other speech-based models by incorporating ASR as an auxiliary task.
+Incorporating ground-truth (GT) personality traits yielded the best performance, with a CCC of 0.766.
+When conditioning on predicted personality labels, our model also surpassed both existing speech- and text-based approaches and achieved comparable performance with multimodal system using both speech and text.
+It is important to note that text input was used only during training for ASR; during inference, the model relied solely on speech.
+Therefore, the most appropriate comparison is with other speech-based systems.
 VII Conclusions
@@
 This work not only addresses the gap in available emotion and personality annotated datasets but also establishes a foundation for future research into multimodal and context-aware affective computing.
+While the personality annotations were derived from acted performances, we also aim to extend our study to more naturalistic data in future work.
 VIII Acknowledgment
-This work was supported by JST SPRING (JPMJSP2110), JST Moonshot R&\&&D (JPMJPS2011).
+This work was supported by JST SPRING (JPMJSP2110), and JST Moonshot R&\&D (JPMJPS2011).
+[Uncaptioned image] Dr.
+Yuan Gao received the B.E.
+degree in school of computer science from Hebei University of Technology, Tianjin, China, in 2019, and the M.S.
+degree from both Tianjin University, Tianjin, China, and the Japan Advanced Institute of Science and Technology, Ishikawa, Japan, in 2022.
+He received the Ph.D.
+degree in Intelligence Science and Technology from Kyoto University, Kyoto, Japan, in 2025.
+He is currently a Research Scientist with SB Intuitions.
+His research interests include speech signal processing and multimodal emotion recognition.
+[Uncaptioned image] Dr.
+Hao Shi (Member, IEEE) received the B.E.
+degree in computer science from Southwest Jiaotong University, Chengdu, China, in 2018, and the M.S.
+degree in computer science from Tianjin University, Tianjin, China, in 2021.
+He received the Ph.D.
+degree in informatics from Kyoto University, Kyoto, Japan, in 2024.
+From October 2024 to March 2025, he was a researcher at Kyoto University.
+He is currently a research scientist with SB Intuitions.
+His research interests include automatic speech recognition and speech enhancement.
+[Uncaptioned image] Dr.
+Yahui Fu received her M.S.
+degrees from both Tianjin University, Tianjin, China, and the Japan Advanced Institute of Science and Technology, Ishikawa, Japan, in 2021.
+She received her Ph.D.
+degree in Intelligence Science and Technology from Kyoto University, Kyoto, Japan, in 2024.
+She is currently a postdoctoral researcher at Kyoto University.
+Her research interests include dialogue systems and multimodal emotion recognition.
+[Uncaptioned image] Chenhui Chu received his B.S.
+in software engineering from Chongqing University in 2008, and his M.S.
+and Ph.D.
+in Informatics from Kyoto University in 2012 and 2015, respectively.
+He is currently a associate professor at Kyoto University.
+His research interests include natural language processing, particularly machine translation and multimodal machine learning.
+[Uncaptioned image] Tatsuya Kawahara (Fellow, IEEE) received B.E.
+in 1987, M.E.
+in 1989, and Ph.D.
+in 1995, all in information science, from Kyoto University, Kyoto, Japan.
+From 1995 to 1996, he was a Visiting Researcher at Bell Laboratories, Murray Hill, NJ, USA.
+Currently, he is a Professor of School of Informatics, Kyoto University.
+From 2020 to 2023, he was the Dean of the School.
+Before that, he was also an Invited Researcher at ATR and NICT.
+He has published more than 450 academic papers on automatic speech recognition, spoken language processing, and spoken dialogue systems.
+He has been conducting several projects including open-source speech recognition software Julius, the automatic transcription system deployed in the Japanese Parliament (Diet), and the autonomous android ERICA.
+Dr.
+Kawahara received the Commendation for Science and Technology by the Minister of Education, Culture, Sports, Science and Technology (MEXT) in 2012.
+From 2003 to 2006, he was a member of IEEE SPS Speech Technical Committee.
+He was a General Chair of IEEE ASRU 2007 and is a General Chair of SIGdial 2024.
+He also served as a Tutorial Chair of INTERSPEECH 2010, a Local Arrangement Chair of ICASSP 2012, and a General Chair of APSIPA ASC 2020.
+He was an editorial board member of Elsevier Journal of Computer Speech and Language and IEEE/ACM Transactions on Audio, Speech, and Language Processing.
+From 2018 to 2021, he was the Editor-in-Chief of APSIPA Transactions on Signal and Information Processing.
+Dr.
+Kawahara is the President of APSIPA, the Secretary General of ISCA, and a Fellow of IEEE.